OpenStudy (anonymous):

Find two numbers whose product is 2 and the sum of whose reciprocals is 33/8.

OpenStudy (pfenn1):

Let x and y be two numbers. The product is 2 means\[xy=2\]The sum of whose reciprocals is 33/8 means\[\frac 1 x +\frac 1y =\left( 33 \over 8 \right)\] Can you solve for x and y?

OpenStudy (anonymous):

Would I start by multiplying the bottom equation by xy?

OpenStudy (pfenn1):

Sure, that is one way to solve.